# Model Merging in LLMs, MLLMs, and Beyond: Methods, Theories, Applications and Opportunities

### v3

@@ 1 Introduction
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [ Model Merging: Methods, Theories, Applications , ver, color=hidden-draw, fill=mygray!100, text width=21.8em, text=black [ Methods (§2), fill=myred!80, text width=8.0em, text=black [ Pre-Merging Mehtods (§2.2), fill=myred!60, text width=14em, text=black [ Linearization Fine-tuning (§2.2.1)
@@
Cross-modal Knowledge Transfer (§3.2.2) , color=hidden-draw, fill=myblue!40, text width=20em, text=black ] ] [ Image Generative Models (§3.3), , fill=myblue!60, text width=17em, text=black [ Style Mixing (§3.3.1)
@@
Second, which applications can benefit from model merging? We discuss in detail the various use cases of model merging in foundation models (§3) and over ten subfields of machine learning (§4). As shown in Figure 2 (lower part), model merging can be applied to a variety of foundation models, including large language models, multimodal large language models, and image generative models. For example, model merging in large language models can help mitigate untruthfulness and toxicity output, accomplish knowledge unlearning, and speed up training. Moreover, model merging also arises in different machine learning subfields, such as continual learning, multi-task/multi-domain learning, few-shot learning, and other subfields, to solve a variety of challenges. For instance, in continual learning, model merging can mitigate catastrophic forgetting of old tasks. In multi-task learning, multi-objective learning and multi-domain learning, it facilitates knowledge transfer. Additionally, in adversarial learning, model merging can be employed for both attack and defense strategies.
@@ 2.3.2 Weighted-based Merging Methods
The above-sophisticated weighting methods operate at the model (or task) level. It is well known that each layer and even each neuron in a deep neural network model play a significantly different role, and some research has developed more fine-grained weighted merging strategies. For example, Layer-wise AdaMerging [194] and aTLAS [205] adaptively learn different sets of merging coefficients for each layer or module of the model, respectively. RegMean [78] indicates that closed-form solutions (relying on the data statistics provided by the training set) exist for linear layers in model merging, while nonlinear layers can simply perform weight averaging. Other works utilize the Fisher information matrix [40] to assess the importance of parameters when merging. Fisher-Merging [113] performs model merging based on the importance of the parameters in each independent model, that is, $\small\Theta^{(\text{merge})}=\sum_{t=1}^{T}F^{(t)}\Theta^{(t)}/\sum_{t=1}^{T}F^{(t)}$, where $F^{(t)}$ is the diagonal of the Fisher information matrix with respect to task $t$. Fisher-nodes-merging [164] also combines a set of Transformer [169] models based on the Fisher information matrix. MaTS [155] developed a block diagonal approximation for Fisher merging. Daheim et al. [29] linked the inaccuracy of weighted average with gradient mismatch, and further proposed an uncertainty-based algorithm to reduce the matching error, ultimately merging the models based on a second-order Hessian estimation.
@@ 2.3.3 Subspace-based Merging Methods
Another class of advanced methods transforms models into sparse subspaces for merging, thereby mitigating task interference. The over-parameterized nature of neural networks and the success of model pruning [54, 22] show that removing most of the parameters from the model barely affects its accuracy [190]. This insight opens up new opportunities for model merging, allowing us to remove insignificant neurons from a single model and merge multiple sparse models within the parameter subspace, as shown in Figure 5 (b).
TIES-Merging [190] proposes to trim each individual model based on parameter magnitudes, retaining only the top 20% of parameters with the highest magnitudes. It further suggests eliminating parameter sign conflicts to reduce interference, and finally merging sparse models using Task Arithmetic [65]. Similarly, Drop And REscale (DARE) [200] also sparsifies by parameter magnitude, and highlights the importance of further performing rescaling on sparse models. In addition to removing the tail parameters with the smallest weight, the Model Breadcrumbs [30] highlight the importance of removing the parameters (outliers) with the largest weights to further reduce noise in model merging and enhance generalization to hyperparameters. TALL-masks [176] creates a mask matrix specific to each task based on a predefined threshold related to independent models. Further, Model Tailor [218] masks unimportant parameters based on the sensitivity of fine-tuned parameters to loss changes and the significance of changes compared to pre-trained parameters. Unlike the standard practice of obtaining a single model through model merging, EMR-Merging [62] proposes maintaining a shared model among multiple tasks alongside a sparse task-specific model. In this approach, the value of the shared model at each index is the largest parameter value among all models. In contrast to the mask construction rules of the aforementioned heuristics, Concrete [156] frames mask construction and model merging as a learnable bi-level optimization problem. The outer-level optimizes the mask matrix, while the inner-level merges the model based on the mask matrix and optimizes it using the unlabeled test samples.
@@ 2.3.4 Routing-based Merging Methods
The basic, weighted-based, or subspace-based merging methods discussed in §2.3.1, §2.3.2 and §2.3.3 are static merging methods. This means that the merged model remains the same for all samples or tasks. Given that there are differences between input samples/tasks, the model’s ability may vary when processing different samples/tasks. As shown in Figure 5 (c), some works propose to dynamically merge models (or subsets of layers) based on the samples/tasks [93, 116, 159, 108] during the inference phase.
@@ 2.4 Theories and Analysis of Model Merging
Second, some studies explain the merging of multiple models with different hyperparameter fine-tuning for the same dataset in terms of connectivity and flatness of the loss landscape. Specifically, some works apply the theory of linear mode connectivity (LMC) [47, 37, 162] of neural networks to explain model merging. LMC reveals that neural network loss minima are not isolated points in the weight space. Recent studies [43, 117, 38, 217] have shown that two independent models, starting from the same pre-trained model and fine-tuned with different configurations, usually satisfy LMC. In other words, LMC is a general phenomenon that typically appears in fine-tuned models based on the ”pretraining-finetuning” paradigm, which is the current standard in the machine learning community. Therefore, performing weight alignment according to LMC provides a robust validity guarantee for model merging [5, 80]. On the other hand, other studies explain model merging from the perspective of a flatter loss landscape [88], arguing that merging multiple weights fine-tuned under different optimization configurations with the same data usually converges to a flat local minimum [41], thus revealing why model merging has better generalization [149, 67, 50, 206, 15].
@@ 3 Application of Model Merging in Foundation Models
Image Generative Models (§3.3) Generate images with multiple styles or achieve image-style transformation
@@ 3.1.1 Human Preference Alignment for LLMs
Some works propose to achieve better, safer, or faster alignment of human preferences by model merging. For example, ExPO [213] adds a task vector, constructed by a moderate model aligned using DPO or RLHF on a small amount of human preference data, to an unaligned SFT model. A more powerful aligned model can be directly obtained by setting a suitable merging coefficient. On the AlpacaEval 2.0 benchmark [97], fusing a model aligned on the 10%/20% preference data with an SFT model results in performance comparable to that of a model aligned on the full preference data. DogeRM [98] proposed merging the reward model with LLMs fine-tuned on different downstream domains to create domain-private reward models directly. Additionally, Lu et al. [107] propose an Online Merging Optimizer, that interpolates the gradient with the SFT model at each step of RLHF. This approach encourages RLHF to optimize toward reward maximization while preventing LLMs from forgetting general knowledge due to RLHF. Beyond preference alignment, several studies have examined the impact of model merging for secure alignment of LLMs [11, 199, 51]. For example, Hammoud et al. [51] find that merging two security-aligned models could compromise security. Thus, they proposed explicitly including secure alignment as an optimization objective when constructing synthetic data for model merging.
@@ 3.1.2 Detoxifcation of LLMs
LLMs have been widely noted for issues related to untruthfulness and toxicity in various applications [60], such as insults, threats, and profanity in responses to certain questions. To address the potential security risks in the application of LLMs, flexible techniques are needed to reduce the generation of toxic text, essentially detoxifying LLMs. A straightforward solution is to collect additional non-toxic data to fine-tune LLMs [83]; however, this approach requires significant computing resources and may interfere with the general capabilities of LLMs. Alternatively, directly reducing the probability of potentially toxic words during the decoding stage requires additional guidance information [87]. Recent studies have shown that reducing the toxic data generation of LLMs through model merging is a simple and effective scheme [65, 209, 60].
@@ 3.1.4 Faster Training of LLMs
Training LLMs requires numerous iterations on massive data, making the training process extremely expensive. For example, training LLAMA2-70B with 2T tokens required 1,720,320 GPU hours [100]. Methods to accelerate LLM training include mixed-precision training, continual retraining, and pipeline parallelism. An orthogonal approach is checkpoint merging in training trajectories, which offers a simple and effective means to either speed up LLM training or enhance training performance at the same cost.
@@ 3.1.5 Combine the Capabilities of Expert LLMs
LLMs exhibit strong generalizability in general tasks, but often lack knowledge in specific vertical domains. Pretrained LLMs typically require fine-tuning within different corporations to become expert LLMs in various fields. Integrating the expertise of multiple specialists is particularly critical for solving more complex tasks. Research on model merging techniques indicates that a composite LLM can be created by combining the parameters of different expert LLMs [200, 6, 31, 215, 171, 172, 201]. For example, Dekoninck et al. [31] demonstrate the ability to flexibly control text generation by merging multiple LLMs with different styles and applying personalized weighting. Robust Weight Signatures [14] proposes a robustness “patching” framework via model merging to enhance the overall robustness of the model against various naturally corrupted versions of clean data. In summary, model merging offers a straightforward and effective strategy for enhancing LLM’s capabilities.
@@ 3.2.2 Model Merging for Cross-Modal Knowledge Transfer
3.3 Model Merging in Image Generative Models
The goal of image generative models, such as generative adversarial networks (GANs), variational autoencoders (VAEs), normalizing flows (Flows), and denoising diffusion probabilistic models (Diffusions), is to approximate the underlying data distribution behind a given dataset so as to generate more new samples with the same distribution. However, image generative models still face the following challenges: the inability to flexibly generate samples with multiple style combinations, the high cost of generative model training, and the inability to generate all the details specified in the instructions. This dilemma has led to an interest in expert models, which train a set of experts with specific abilities on different data shards or distributions, allowing for the flexible addition or removal of certain styles of experts at inference time. Considering the difficulty in deploying and the cost of resources for ensemble learning, model merging offers a new perspective on combining skill-specific experts of different styles without additional memory and inference costs.
@@ 4.1.1 Model Merging to Mitigate Catastrophic Forgetting
Tangent Model Composition [104] proposes fine-tuning each task independently in the tangent space of the pre-trained model and then linearly fine-tuning these models to perform CL. This approach does not depend on the specific settings of CL and can be easily applied to task, class, and domain-incremental learning scenarios. In addition, ITA [127] emphasizes the necessity for the fine-tuned model to be in the same basin as the pre-trained model to ensure the composability of nonlinear models. It introduces a regularization term similar to EWC [85] in traditional CL to constrain the distance between the fine-tuned weights and the pre-trained weights when training the independent model. WARP [135] suggests linearly interpolating the pre-trained LLM’s weights with its aligned weights via RLHF on a preference dataset, thus mitigating the forgetting of knowledge from the pre-trained LLM. BAM [7] continuously adapts LLMs to new languages by merging models while preserving general capabilities. Model Tailor [218] explores the problem of catastrophic forgetting during fine-tuning of MLLMs, and proposes to merge only the most important subset of parameters in the fine-tuned MLLM model into the pre-trained MLLM model, so as to retain the generalization ability of the pre-trained model as much as possible, while compensating the selected weights to reduce the performance of the fine-tuning task. MagMax [112] merges pruned task vectors to further alleviate parameter sign conflicts and old knowledge forgetting. Equifinality, PAINT [64] and LM-Cocktail [187] interpolate the weights of the fine-tuned model and the zero-shot model to improve accuracy on downstream tasks without degrading accuracy on supported/general tasks.
@@ 4.2.1 Knowledge Transfer in Multi-Task Learning
The goal of multi-task learning (MTL) is to enable a single model to perform multiple tasks simultaneously, thereby facilitating knowledge transfer between these tasks [110, 18, 144, 152, 202, 192]. As shown in Figure 1(c), to avoid the high cost of joint training, a straightforward approach is to merge multiple independently trained models on different tasks to accomplish MTL. Almost all of the model merging methods discussed in §2.3 can be used to merge multiple models trained on different tasks to perform MTL. In this section, we take some representative tasks as examples. For MTL tasks in computer vision, Task Arithmetic [65], Ties-Merging [190], AdaMerging [194] and other studies [197, 156, 193] proposed to combine ViT models trained on different visual classification tasks, and the obtained model can complete the object classification of multiple tasks. The results of Task Arithmetic [65] demonstrate that merging independently trained models from any two datasets yields a merged model whose performance is comparable to that of a single-task model. Similarly, ZipIt [151], which merges ResNet architectures trained on different tasks, achieves comparable results. For MTL tasks in natural language processing, DARE [200] introduces a method to assimilate homologous models, augmenting LLMs as a ”free lunch”. For instance, merging WizardLM with WizardMath significantly boosts WizardLM’s performance on GSM8K (a benchmark for evaluating the mathematical reasoning ability of LLMs) from 2.2 to 66.3. Akiba et al. [6] suggest that directly merging an LLM with mathematical capabilities and an LLM with Japanese language proficiency results in a model capable of solving Japanese mathematical problems. Furthermore, numerous studies have demonstrated that combining PEFT modules (such as Adapter or LoRA) trained on different tasks can also achieve MTL [160, 208].

### v4

@@ 1 Introduction
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [ Model Merging: Methods, Theories, Applications , ver, color=hidden-draw, fill=mygray!100, text width=19.5em, text=black [ Methods (§2), fill=myred!80, text width=8.0em, text=black [ Pre-Merging Mehtods (§2.2), fill=myred!60, text width=14em, text=black [ Linearization Fine-tuning (§2.2.1)
@@
Cross-modal Knowledge Transfer (§3.2.2) , color=hidden-draw, fill=myblue!40, text width=20em, text=black ] ] [ Visual Generative Models (§3.3), , fill=myblue!60, text width=17em, text=black [ Style Mixing (§3.3.1)
@@
Second, which applications can benefit from model merging? We discuss in detail the various use cases of model merging in foundation models (§3) and over ten subfields of machine learning (§4). As shown in Figure 2 (lower part), model merging can be applied to a variety of foundation models, including large language models, multimodal large language models, and visual generative models. For example, model merging in large language models can help mitigate untruthfulness and toxicity output, accomplish knowledge unlearning, and speed up training. Moreover, model merging also arises in different machine learning subfields, such as continual learning, multi-task/multi-domain learning, few-shot learning, and other subfields, to solve a variety of challenges. For instance, in continual learning, model merging can mitigate catastrophic forgetting of old tasks. In multi-task learning, multi-objective learning and multi-domain learning, it facilitates knowledge transfer. Additionally, in adversarial learning, model merging can be employed for both attack and defense strategies.
@@ 2.3.2 Weighted-based Merging Methods
The above-sophisticated weighting methods operate at the model (or task) level. It is well known that each layer and even each neuron in a deep neural network model play a significantly different role, and some research has developed more fine-grained weighted merging strategies. For example, Layer-wise AdaMerging [220] and aTLAS [231] adaptively learn different sets of merging coefficients for each layer or module of the model, respectively. RegMean [86] indicates that closed-form solutions (relying on the data statistics provided by the training set) exist for linear layers in model merging, while nonlinear layers can simply perform weight averaging. Other works utilize the Fisher information matrix [44] to assess the importance of parameters when merging [126, 80]. Fisher-Merging [126] performs model merging based on the importance of the parameters in each independent model, that is, $\small\Theta^{(\text{merge})}=\sum_{t=1}^{T}F^{(t)}\Theta^{(t)}/\sum_{t=1}^{T}F^{(t)}$, where $F^{(t)}$ is the diagonal of the Fisher information matrix with respect to task $t$. Fisher-nodes-merging [186] also combines a set of Transformer [191] models based on the Fisher information matrix. MaTS [175] developed a block diagonal approximation for Fisher merging. Daheim et al. [31] linked the inaccuracy of weighted average with gradient mismatch, and further proposed an uncertainty-based algorithm to reduce the matching error, ultimately merging the models based on a second-order Hessian estimation.
@@ 2.3.3 Subspace-based Merging Methods
Another class of advanced methods transforms models into sparse subspaces for merging, thereby mitigating task interference. The over-parameterized nature of neural networks and the success of model pruning [59, 24] show that removing most of the parameters from the model barely affects its accuracy [216]. This insight opens up new opportunities for model merging, allowing us to remove insignificant neurons from a single model and merge multiple sparse models within the parameter subspace [33, 60], as shown in Figure 5 (b).
TIES-Merging [216] proposes to trim each individual model based on parameter magnitudes, retaining only the top 20% of parameters with the highest magnitudes. It further suggests eliminating parameter sign conflicts to reduce interference, and finally merging sparse models using Task Arithmetic [71]. Similarly, Drop And REscale (DARE) [226] also sparsifies by parameter magnitude, and highlights the importance of further performing rescaling on sparse models. In addition to removing the tail parameters with the smallest weight, the Model Breadcrumbs [32] highlight the importance of removing the parameters (outliers) with the largest weights to further reduce noise in model merging and enhance generalization to hyperparameters. TALL-masks [200] creates a mask matrix specific to each task based on a predefined threshold related to independent models. Further, Model Tailor [243] masks unimportant parameters based on the sensitivity of fine-tuned parameters to loss changes and the significance of changes compared to pre-trained parameters. APL [96] proposes estimating parameter importance using causal interventions. Unlike the standard practice of obtaining a single model through model merging, EMR-Merging [68] proposes maintaining a shared model among multiple tasks alongside a sparse task-specific model. In this approach, the value of the shared model at each index is the largest parameter value among all models. In contrast to the mask construction rules of the aforementioned heuristics, Concrete [176] frames mask construction and model merging as a learnable bi-level optimization problem. The outer-level optimizes the mask matrix, while the inner-level merges the model based on the mask matrix and optimizes it using the unlabeled test samples.
@@ 2.3.4 Routing-based Merging Methods
The basic, weighted-based, or subspace-based merging methods discussed in §2.3.1, §2.3.2 and §2.3.3 are static merging methods. This means that the merged model remains the same for all samples or tasks. Given that there are differences between input samples/tasks, the model’s ability may vary when processing different samples/tasks. As shown in Figure 5 (c), some works propose to dynamically merge models (or subsets of layers) based on the samples/tasks [105, 130, 179, 121, 91] during the inference phase.
@@ 2.4 Theories and Analysis of Model Merging
Second, some studies explain the merging of multiple models with different hyperparameter fine-tuning for the same dataset in terms of connectivity and flatness of the loss landscape [98, 166, 202, 12]. Specifically, some works apply the theory of linear mode connectivity (LMC) [51, 41, 183] of neural networks to explain model merging. LMC reveals that neural network loss minima are not isolated points in the weight space. Recent studies [47, 131, 42, 242] have shown that two independent models, starting from the same pre-trained model and fine-tuned with different configurations, usually satisfy LMC. In other words, LMC is a general phenomenon that typically appears in fine-tuned models based on the ”pretraining-finetuning” paradigm, which is the current standard in the machine learning community. Therefore, performing weight alignment according to LMC provides a robust validity guarantee for model merging [5, 88]. On the other hand, other studies explain model merging from the perspective of a flatter loss landscape [100], arguing that merging multiple weights fine-tuned under different optimization configurations with the same data usually converges to a flat local minimum [45], thus revealing why model merging has better generalization [168, 73, 54, 232, 17].
@@ 3 Application of Model Merging in Foundation Models
Visual Generative Models (§3.3) Generate images with multiple styles or achieve image-style transformation
@@ 3.1.1 Human Preference Alignment for LLMs
Some works propose to achieve better, safer, or faster alignment of human preferences by model merging [238, 199, 13, 48, 141]. For example, ExPO [238] adds a task vector, constructed by a moderate model aligned using DPO or RLHF on a small amount of human preference data, to an unaligned SFT model. A more powerful aligned model can be directly obtained by setting a suitable merging coefficient. On the AlpacaEval 2.0 benchmark [110], fusing a model aligned on the 10%/20% preference data with an SFT model results in performance comparable to that of a model aligned on the full preference data. DogeRM [111] proposed merging the reward model with LLMs fine-tuned on different downstream domains to create domain-private reward models directly. Additionally, Lu et al. [120] propose an Online Merging Optimizer, that interpolates the gradient with the SFT model at each step of RLHF. This approach encourages RLHF to optimize toward reward maximization while preventing LLMs from forgetting general knowledge due to RLHF. Beyond preference alignment, several studies have examined the impact of model merging for secure alignment of LLMs [13, 225, 55, 57]. For example, Hammoud et al. [55] find that merging two security-aligned models could compromise security. Thus, they proposed explicitly including secure alignment as an optimization objective when constructing synthetic data for model merging.
@@ 3.1.2 Detoxifcation of LLMs
LLMs have been widely noted for issues related to untruthfulness and toxicity in various applications [66], such as insults, threats, and profanity in responses to certain questions. To address the potential security risks in the application of LLMs, flexible techniques are needed to reduce the generation of toxic text, essentially detoxifying LLMs. A straightforward solution is to collect additional non-toxic data to fine-tune LLMs [92]; however, this approach requires significant computing resources and may interfere with the general capabilities of LLMs. Alternatively, directly reducing the probability of potentially toxic words during the decoding stage requires additional guidance information [97]. Recent studies have shown that reducing the toxic data generation of LLMs through model merging is a simple and effective scheme [71, 234, 66, 37].
@@ 3.1.4 Faster Training of LLMs
Training LLMs requires numerous iterations on massive data, making the training process extremely expensive [89, 148]. For example, training LLAMA2-70B with 2T tokens required 1,720,320 GPU hours [113]. Methods to accelerate LLM training include mixed-precision training, continual retraining, and pipeline parallelism. An orthogonal approach is checkpoint merging in training trajectories, which offers a simple and effective means to either speed up LLM training or enhance training performance at the same cost.
@@ 3.1.5 Combine the Capabilities of Expert LLMs
LLMs exhibit strong generalizability in general tasks, but often lack knowledge in specific vertical domains. Pretrained LLMs typically require fine-tuning within different corporations to become expert LLMs in various fields. Integrating the expertise of multiple specialists is particularly critical for solving more complex tasks. Research on model merging techniques indicates that a composite LLM can be created by combining the parameters of different expert LLMs [226, 6, 34, 240, 193, 194, 227, 171, 227, 99]. For example, Dekoninck et al. [34] demonstrate the ability to flexibly control text generation by merging multiple LLMs with different styles and applying personalized weighting. Robust Weight Signatures [16] proposes a robustness “patching” framework via model merging to enhance the overall robustness of the model against various naturally corrupted versions of clean data. In summary, model merging offers a straightforward and effective strategy for enhancing LLM’s capabilities.
@@ 3.2.2 Model Merging for Cross-Modal Knowledge Transfer
3.3 Model Merging in Visual Generative Models
The goal of visual generative models, such as generative adversarial networks (GANs), variational autoencoders (VAEs), normalizing flows (Flows), and denoising diffusion probabilistic models (Diffusions), is to approximate the underlying data distribution behind a given dataset so as to generate more new samples with the same distribution. However, image generative models still face the following challenges: the inability to flexibly generate samples with multiple style combinations, the high cost of generative model training, and the inability to generate all the details specified in the instructions. This dilemma has led to an interest in expert models, which train a set of experts with specific abilities on different data shards or distributions, allowing for the flexible addition or removal of certain styles of experts at inference time. Considering the difficulty in deploying and the cost of resources for ensemble learning, model merging offers a new perspective on combining skill-specific experts of different styles without additional memory and inference costs.
@@ 4.1.1 Model Merging to Mitigate Catastrophic Forgetting
To overcome catastrophic forgetting and maintain new learning ability, many works have proposed to merge models for old tasks and update models for new tasks [209, 161, 138]. Tangent Model Composition [117] proposes fine-tuning each task independently in the tangent space of the pre-trained model and then linearly fine-tuning these models to perform CL. This approach does not depend on the specific settings of CL and can be easily applied to task, class, and domain-incremental learning scenarios. In addition, ITA [143] emphasizes the necessity for the fine-tuned model to be in the same basin as the pre-trained model to ensure the composability of nonlinear models. It introduces a regularization term similar to EWC [94] in traditional CL to constrain the distance between the fine-tuned weights and the pre-trained weights when training the independent model. WARP [152] suggests linearly interpolating the pre-trained LLM’s weights with its aligned weights via RLHF on a preference dataset, thus mitigating the forgetting of knowledge from the pre-trained LLM. BAM [7] continuously adapts LLMs to new languages by merging models while preserving general capabilities. Model Tailor [243] explores the problem of catastrophic forgetting during fine-tuning of MLLMs, and proposes to merge only the most important subset of parameters in the fine-tuned MLLM model into the pre-trained MLLM model, so as to retain the generalization ability of the pre-trained model as much as possible, while compensating the selected weights to reduce the performance of the fine-tuning task. MagMax [125] merges pruned task vectors to further alleviate parameter sign conflicts and old knowledge forgetting. Equifinality, PAINT [70] and LM-Cocktail [213] interpolate the weights of the fine-tuned model and the zero-shot model to improve accuracy on downstream tasks without degrading accuracy on supported/general tasks.
@@ 4.2.1 Knowledge Transfer in Multi-Task Learning
The goal of multi-task learning (MTL) is to enable a single model to perform multiple tasks simultaneously, thereby facilitating knowledge transfer between these tasks [123, 20, 162, 172, 228, 218]. As shown in Figure 1(c), to avoid the high cost of joint training, a straightforward approach is to merge multiple independently trained models on different tasks to accomplish MTL. Almost all of the model merging methods discussed in §2.3 can be used to merge multiple models trained on different tasks to perform MTL. In this section, we take some representative tasks as examples. For MTL tasks in computer vision, Task Arithmetic [71], Ties-Merging [216], AdaMerging [220] and other studies [223, 176, 219] proposed to combine ViT models trained on different visual classification tasks, and the obtained model can complete the object classification of multiple tasks. The results of Task Arithmetic [71] demonstrate that merging independently trained models from any two datasets yields a merged model whose performance is comparable to that of a single-task model. Similarly, ZipIt [170], which merges ResNet architectures trained on different tasks, achieves comparable results. For MTL tasks in natural language processing, DARE [226] introduces a method to assimilate homologous models, augmenting LLMs as a ”free lunch”. For instance, merging WizardLM with WizardMath significantly boosts WizardLM’s performance on GSM8K (a benchmark for evaluating the mathematical reasoning ability of LLMs) from 2.2 to 66.3. Akiba et al. [6] suggest that directly merging an LLM with mathematical capabilities and an LLM with Japanese language proficiency results in a model capable of solving Japanese mathematical problems. Furthermore, numerous studies have demonstrated that combining PEFT modules (such as Adapter, LoRA or Soft-prompts) trained on different tasks can also achieve MTL [180, 234, 11].